OpenStudy (anonymous):

Multiplying Radicals. 48x^5y^3/y^4* x^2y/6x^3y^2

OpenStudy (turingtest):

I don't see any radicals, but you should know here that\[x^a\cdot x^b=x^{a+b}\]and\[\frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (anonymous):

whoops, no radicals, just expressions. sorry.

OpenStudy (akshay_budhkar):

\[48x^5y^3/y^4* x^2y/6x^3y^2\] seems to be your question.. just use turing's formulas to get the answer

OpenStudy (anonymous):

im stuck on the factoring part.

OpenStudy (anonymous):

hello? :(

jhonyy9 (jhonyy9):

48x^5y^3 x^2y 8x^4 ---------- * --------- = ----- y^4 6x^3y^2 y^2

jhonyy9 (jhonyy9):

48x^5y^3 x^2y 8x^4 ---------- * --------- = ----- y^4 6x^3y^2 y^2

jhonyy9 (jhonyy9):

48x^5y^3 x^2y 8x^4 ---------- * --------- = ----- y^4 6x^3y^2 y^2